OpenStudy (anonymous):

Can anyone explain this to me? Thanks "Let f(x)=-2+4 and g(x)=-6x-7. Find f(x)-g(x)."

OpenStudy (solomonzelman):

\(\large\color{black}{ f(x)=-2x+4 }\) ( this is the correct, f(x), right? ) \(\large\color{black}{ g(x)=-6x-7 }\)

OpenStudy (solomonzelman):

\(\large\color{red}{ f(x)=-2x+4 }\) \(\large\color{blue}{ g(x)=-6x-7 }\) So, \(\large\color{black}{ \color{red}{ f(x)}~-~\color{blue}{g(x)} =\color{red}{(-2x+4)}~-~\color{blue}{(-6x-7)} }\)

OpenStudy (skullpatrol):

$$\Huge a-b=a+(-b)$$

OpenStudy (skullpatrol):

$$\Huge -(-a)=a$$

OpenStudy (solomonzelman):

And, \(\normalsize\color{ slate }{\Huge{\bbox[5pt, lightcyan ,border:2px solid blue ]{ -c(-a-b)=ca+cb}}}\)

OpenStudy (skullpatrol):

Any questions @sammiekate1 ?

OpenStudy (anonymous):

That makes sense, thank you all very much!!

OpenStudy (skullpatrol):

Thanks for asking :)